# Fibred surfaces and their unitary rank

### v1

@@ 4 Some new slope inequalities
We prove in this section several slope inequalities involving the unitary rank $u_{f}$, and some also involving $c_{f}$, and the rank $r_{1}$ of the first subbundle $\mathcal{E}_{1}$.
@@ 4.1 First applications
Even avoiding the elaboration made in Lemmas 3.16 and 3.8 one can derive directly from Proposition 3.13 some interesting slope inequalities, which will be very important for the applications of the following sections.
@@
Let $f\colon S\to B$ a relatively minimal non-locally trivial fibred surface of genus $g\geq 2$. Let $\alpha:=\alpha_{1}$ be the maximal multiplicity of the non-vertical irreducible components of $Z_{1}$, the following inequality holds:
@@
Moreover, equality holds in (4.3) if and only if $u_{f}=g-1$ and $K_{f}$ is numerically equivalent to $(2g-2)\Gamma+\chi_{f}F=(2g-2)\Gamma+\mu_{1}F$, where $\Gamma$ is a section of $f$.
@@
First observe that the assumption that $f$ is non-locally trivial implies that $\chi_{f}>0$ and so equivalently (Remark 3.5) $u_{f}\leq g-1$.
From Konno’s result (Proposition 3.13) we have that $(\alpha+1)K_{f}-\mu_{1}F$ is a nef divisor on $S$, and by Xiao’s result (Proposition 3.9) $K_{f}-\mu_{1}F\equiv M_{1}-\mu_{1}F+Z_{1}$ is a pseudoeffective divisor, so their intersection is greater or equal to 00:
$((\alpha+1)K_{f}-\mu_{1}F\cdot K_{f}-\mu_{1}F)=(\alpha+1)K_{f}^{2}-(2g-2)(\alpha+2)\mu_{1}\geq 0,$
@@
$K_{f}^{2}\geq(2g-2)\frac{\alpha+2}{\alpha+1}\mu_{1}\geq\frac{(\alpha+2)(2g-2)}{(\alpha+1)(g-u_{f})}\chi_{f}.$
So, inequality (4.2) is proved.
Now just observe that the function
@@
is strictly decreasing and $\alpha\leq 2g-2$ so we obtain inequality (4.3).
As for the case where equality holds, we see that we need to have equalities in all the passages of the proof, so in particular $\alpha=a=2g-2$, i.e $d_{1}=0$, so necessarily we have (see Remark 3.7) $r_{1}=1$. Moreover, we need to have $\mu_{1}=\chi_{f}/(g-u_{f})=\mu(\mathcal{A})$, but this only can happen if $\mathcal{A}$ is semistable (see $(ii)$ of Remark 3.6) hence $l=2$. Putting all together, we have that necessarily $g-u_{f}=\mbox{{rank}}\mathcal{A}=\mbox{{rank}}\mathcal{E}_{1}=1$ and $\mu_{1}=\chi_{f}$. Note moreover that, as $\alpha=a=2g-2$ the divisor $Z_{1}$ need to be $(2g-2)\Gamma+Z^{\prime}$, where $\Gamma$ is a section of $f$, and $Z^{\prime}$ is a vertical divisor.
Now we prove that necessarily if equality holds in (4.3) then $Z^{\prime}=0$.
First of all we observe that $Z^{\prime}$ does not contain any fibre in its support (see also Remark 3.10): if this was the case, we would have
$K_{f}\equiv(2g-2)\Gamma+bF+Z^{\prime\prime}+\mu_{1}F=(2g-2)\Gamma+Z^{\prime\prime}+(\mu_{1}+b)F,$
with $b\geq 1$ and $Z^{\prime\prime}\geq 0$ vertical, so $K_{f}-(\mu_{1}+b)F$ would be effective and we would have a sharper inequality than (4.3).
We now prove that there are no vertical components. Suppose on the contrary $Z^{\prime}\not=0$, and let $D$ be an irreducible component of $Z^{\prime}$ (which is not a fibre by the previous argument). We have equality in (4.3)
@@
On the other side $K_{f}^{2}=\frac{4g(g-1)}{2g-1}\mu_{1}+4(g-1)\Gamma\cdot Z^{\prime}$, so we need to have $\Gamma\cdot Z^{\prime}=0$ which implies $K_{f}\cdot Z^{\prime}=Z^{\prime 2}$. But $Z^{\prime 2}\leq 0$ because $Z^{\prime}$ is vertical, and so $K_{f}\cdot Z^{\prime}=Z^{\prime 2}=0$. But the algebraic index theorem [4, Thm. 4.14], as $K_{f}^{2}>0$ (Theorem 2.2) implies that $Z^{\prime 2}<0$assuming $Z^{\prime}\not=0$- which gives the desired contradiction.
@@
On the other side, suppose that $u_{f}=g-1$ and $K_{f}$ is numerically equivalent to $(2g-2)\Gamma+\chi_{f}F$, with $\Gamma$ a section of $f$. Then $K_{f}^{2}=2(2g-2)\chi_{f}+4(g-1)^{2}\Gamma^{2}$. Now, observe that $K_{f}\cdot\Gamma=-\Gamma^{2}$ by adjunction, and that -by the Algebraic index theorem again- the intersection matrix between $K_{f},\Gamma$ and $F$ has zero determinant:
@@
Inequality (4.3) is a a generalization of [34, Lemma 2.7]. For $u_{f}\ll g$ this result is not very interesting, for instance in case $u_{f}=0$ (4.3) gives as a coefficient $(4(g-1))/(2g-1)$, which is strictly smaller than the one of the slope inequality. But for $u_{f}$ large with respect to $g$ both inequalities become interesting. Indeed we will see in Section 6 that the last one implies a Xiao type inequality on $u_{f}$.
We now see how we can use again the Hodge index theorem already mentioned in the proof of Theorem 4.1 in combination with Konno’s method, to obtain a somewhat different inequality as follows. Let $\Sigma$ be the irreducible horizontal component of $Z_{1}$ of maximal multiplicity $\alpha_{=}\alpha_{1}$. If there is more than one such components we simply choose one. Then the restriction of $f$ to $\Sigma$ is a (ramified) covering of $\pi\colon\Sigma\to B$, of degree say $\beta$.
@@
First of all observe that $K_{f}\cdot\Sigma+\Sigma^{2}=r\pi\geq 0$. As $K_{f}^{2}>0$ we have that the intersection matrix between $K_{f},\Sigma$ and $F$ has determinant greater or equal to 00:
@@
Now, plugging this inequality in the argument of Lemma 3.18, and in particular in (3.21), we obtain:
$K_{f}^{2}\geq(4g-4-a)\mu_{1}+K_{f}\cdot Z_{1}\geq(4g-4-a)\mu_{1}+\frac{\alpha\beta^{2}}{4(g-1)(\beta+g-1)}K_{f}^{2};$
and so we obtain the desired inequality. ∎
@@
Observe that in case $a=\alpha=2g-2$ (and thus $\beta=1$), we obtain (4.3). In general inequality (4.6) seems to be weaker than (4.2), unless one has specific informations on $\beta$ and $\alpha$.
@@
Let $f\colon S\to B$ a relatively minimal fibred surface of genus $g\geq 2$. Assume that $Z_{1}$ doesn’t have horizontal components, i.e. $a=\alpha=0$ (equivalently the linear system induced by the stalk of $\mathcal{E}_{1}$ on the general fibre is base point free); Then the following slope inequality holds:
@@
Inequality (4.9) can be deduced from Theorem 4.1 setting $\alpha=0$ in (4.3). ∎
@@
Notice however that inequality (4.9) is false in general, as proved by the examples of Lu in [41]. Moreover, the assumption on the base locus of the system induced by $\mathcal{E}_{1}$ is quite restrictive, and not very easy to verify in concrete cases.
@@
The previous theorem is somewhat close to one due to the author with Barja in [7, Remark 3.3]: there we proved that inequality (4.9) holds if the ample summand $\mathcal{A}$ of the Hodge bundle is semistable and its stalk induces a base point free system on the general fibres. Recall that $\mathcal{A}$ is either the whole $f_{*}\omega_{f}$ or the one to last subsheaf of the Harder-Narashiman sequence $\mathcal{E}_{l}$. So, the assumptions of [7, Remark 3.3] can be rephrased as follows: either the Hodge bundle is semistable (in this case by definition $Z_{1}=0$) or it has a Harder-Narashiman sequence of lenght two, $\mu_{2}=0$ and $Z_{1}$ has no horizontal components. So, the above Theorem implies [7, Remark 3.3].
@@
Let $f\colon S\to B$ a relatively minimal fibred surface of genus $g\geq 2$. Let us denote as usual $r_{1}=\mbox{{rank}}\mathcal{E}_{1}$.
@@
Moreover, suppose that equality is reached. Then:
  • if $r_{1}=1$ then $f$ is a trigonal fibration;
  • if $r_{1}\geq 2$ then $F$ is a hyperelliptic fibration.
@@
If $u_{f}>0$, then by Remark 3.5 we have $\mathcal{A}=\mathcal{E}_{l-1}$, and the Harder-Narashiman sequence of $\mathcal{A}$ is just the truncation of the one of $\mathcal{E}$:
@@
Let us recall inequality (3.8):
$K_{f}^{2}\geq\left(\frac{(\alpha+1)(4g-4-a)+a}{\alpha+1}\right)\mu_{1}.$
Now, observe that $a\geq\alpha$ and that the function
@@
If $r_{1}=1$, then $a=2g-2$, and we obtain (4.3), which coincides with inequality (4.13).
If $r_{1}>1$, then we can apply Clifford’s theorem ([3, Chap.III Sec 1]) to the subcanonical linear system ${M_{i}}_{|F}$ and obtain $d_{1}\geq 2(r_{1}-1)$, so $a_{1}=2g-2-d_{1}\leq 2g-2r_{1}$, so we can use the inequality as follows:
$\begin{split}K_{f}^{2}&\geq\mu_{1}f(2g-2r_{1})=\frac{(2g-2r_{1}+1)(2g+2r_{1}-4)+2g-2r_{1}}{2g-2r_{1}+1}\mu_{1}\geq\\ &\geq\frac{(2g+2r_{1}-4)(2g-2r_{1}+1)+2g-2r_{1}}{2g-2r_{1}+1}\frac{\chi_{f}}{g-u_{f}},\end{split}$
as wanted. Let us now suppose that equality holds in (4.13). If $r_{1}=1$, we proved in Theorem 4.1 that necessarily $u_{f}=g-1$, and so $f$ is a trigonal fibration by Proposition 6.4 below. If $r_{2}\geq 2$ then the fibre of $\mathcal{E}_{1}$ induce a linear series on $F$, of degree necessarily $d_{1}=2r_{1}-2$. So, if $r_{1}<g$, again Clifford’s Theorem tells us that $F$ is hyperelliptic. If $r_{1}=g$, then the Hodge bundle is semistable, and in particular $u_{f}=0$. So, the bound becomes the classical slope inequality (2.15), and we can use Theorem 2.14. ∎
@@
Notice that in the worst case possible for (1) i.e. for $r_{1}=1$, we obtain again precisely the bound (4.3) of Theorem 4.1.
We can somehow improve this inequality by using a sharpening of Clifford’s theorem obtained by the author with E. Riva in [48, Theorem 2.13]. Let us recall the result in the form needed here:
@@
First two simple remarks. If the fibration is hyperelliptic i.e. $c_{f}=0$, we have $m=0$ and so inequality (4.17) becomes precisely inequality (4.13) of the theorem above. If $r_{1}=1$, then inequality (4.17) again becomes (4.3).
So we have reduced ourselves to the case when $F$ is non-hyperelliptic and $r_{1}\geq 2$. In this case we are in the position to apply Theorem 4.15, and the theorem follows immediately by applying Lemma 3.18 observing that
@@
Remark 4.18.
An interesting application of this result is the consequences we can derive in the case that $\mathcal{A}$ is semistable: see the next section.
In case $r_{1}\geq 2$ we can give sharper inequalities as follows:
@@
Then one of the following possibilities hold:
@@
(i) If $h^{1}(R_{1})\geq 2$, then $R_{1}$ contributes to the Clifford index (REF) of $F$ and so $d_{1}\geq 2(h^{0}(R_{1})-1)+\mbox{{Cliff}}(F)\geq 2r_{1}-2+c_{f}$. So we have that $a_{1}\leq 2g-2r_{1}-c_{f}$, and we obtain equation (4.20) by using (3.19).
(ii) If $h^{1}(R_{1})=1$, in theory $R_{1}$ could be 00, but in this case we would have necessarily $d_{1}=0$ and so $r_{1}=1$, contrary to the assumption. So, necessarily $R_{1}\not=0$, and by Riemann-Roch
@@
hence $a_{1}\leq g-r_{1}$, and we just use this bound again in (3.19). ∎
4.3 Slope inequalities for Harder-Narashiman sequence of lenght $\leq 2$
Now we prove new slope inequalities for the case where the Harder-Narashiman sequence is short. For the case $\ell=1$, i.e $f_{*}\omega_{f}$ is semistable, recall the following beautiful result:
Theorem 4.22 (Lu-Zuo [39]).

### v2

@@ 4 Some new slope inequalities
We prove in this section several slope inequalities involving the unitary rank $u_{f}$, and some also involving $c_{f}$, and the rank $r_{1}$ of the first subbundle $\mathcal{E}_{1}$ of the Harder-Narasimhan filtration of the Hodge bundle.
@@ 4.1 First applications
Even avoiding the elaboration made in Lemmas 3.15 and 3.7, one can derive directly from Proposition 3.12 some interesting slope inequalities, which will be very important for the applications in the following sections.
@@
Let $f\colon S\to B$ be a relatively minimal locally non-trivial fibred surface of genus $g\geq 2$. Let $\alpha:=\alpha_{1}$ be the maximal multiplicity of the non-vertical irreducible components of $Z_{1}$. The following inequality holds:
@@
Moreover, the equality holds in (4.3) if and only if $u_{f}=g-1$ and $K_{f}$ is numerically equivalent to $(2g-2)\Gamma+\chi_{f}F=(2g-2)\Gamma+\mu_{1}F$, where $\Gamma$ is a section of $f$.
@@
First observe that the assumption that $f$ is locally non-trivial implies that $\chi_{f}>0$ and so equivalently (Remark 3.4) $u_{f}\leq g-1$. From Proposition 3.12 we have that $(\alpha+1)K_{f}-\mu_{1}F$ is a nef divisor on $S$, and by Proposition 3.8 $K_{f}-\mu_{1}F\equiv M_{1}-\mu_{1}F+Z_{1}$ is a pseudo-effective divisor, so their intersection is greater than or equal to 00:
$((\alpha+1)K_{f}-\mu_{1}F)\cdot(K_{f}-\mu_{1}F)=(\alpha+1)K_{f}^{2}-(2g-2)(\alpha+2)\mu_{1}\geq 0.$
@@
$K_{f}^{2}\geq(2g-2)\frac{\alpha+2}{\alpha+1}\mu_{1}.$
Therefore, the inequality (4.2) is proved. Now just observe that the function
@@
is strictly decreasing and $\alpha\leq 2g-2$ so we obtain
$K_{f}^{2}\geq\frac{(2g-2)(2g)}{2g-1}\mu_{1}\geq\frac{(2g-2)(2g)}{2g-1}\frac{\chi_{f}}{g-u_{f}},$
where the last inequality follows from Remark 3.5.
Suppose now that the equality holds in (4.3). Then the equalities also hold in all the passages of the proof. In particular:
  • $\alpha=a=2g-2$, i.e $d_{1}=0$, so by Remark 3.6 we have that $r_{1}=1$;
  • $\mu_{1}=\chi_{f}/(g-u_{f})=\mu(\mathcal{A})$, so by $(ii)$ of Remark 3.5 $\mathcal{A}$ is semistable.
Putting all together, we have that necessarily the Harder-Narasimhan filtration has lenght $2$, that $g-u_{f}=\mbox{{rank}}\mathcal{A}=\mbox{{rank}}\mathcal{E}_{1}=1$ and that $\mu_{1}=\chi_{f}$.
The fact that $\alpha=a=2g-2$ implies by Remark 3.14 that the divisor $Z_{1}$ is necessarily $(2g-2)\Gamma+Z^{\prime}$, where $\Gamma$ is a section of $f$, and $Z^{\prime}$ is an effective vertical divisor.
Now we prove that if the equality holds in (4.3) then $Z^{\prime}=0$.
First of all observe that, via the same arguments used in Remark 3.9, we see that $Z^{\prime}$ does not contain any fibre in its support
We now prove that there are no vertical components. Suppose on the contrary $Z^{\prime}\not=0$, and let $D$ be an irreducible component of $Z^{\prime}$. We have the equality in (4.3)
@@
On the other side $K_{f}^{2}=\frac{4g(g-1)}{2g-1}\mu_{1}+4(g-1)\Gamma\cdot Z^{\prime}$, so we need to have $\Gamma\cdot Z^{\prime}=0$ which implies $K_{f}\cdot Z^{\prime}=Z^{\prime 2}$. But $Z^{\prime 2}\leq 0$ because $Z^{\prime}$ is vertical, and so $K_{f}\cdot Z^{\prime}=Z^{\prime 2}=0$. The algebraic index theorem [2, Thm. 4.14], gives a contradiction: $K_{f}^{2}>0$ implies that $Z^{\prime 2}<0$.
@@
Now we prove the converse implication. Suppose that $u_{f}=g-1$ and $K_{f}$ is numerically equivalent to $(2g-2)\Gamma+\chi_{f}F$, with $\Gamma$ a section of $f$. Then $K_{f}^{2}=2(2g-2)\chi_{f}+4(g-1)^{2}\Gamma^{2}$. Now, observe that $K_{f}\cdot\Gamma=-\Gamma^{2}$ by adjunction, and that –by the algebraic index theorem again– the intersection matrix between $K_{f},\Gamma$ and $F$ has zero determinant:
@@
The inequality (4.3) is a generalization of [21, Lemma 2.7]. For $u_{f}\ll g$ this result is not very interesting; for example, in case $u_{f}=0$ (4.3) gives as a coefficient $(4(g-1))/(2g-1)$, which is strictly smaller than the one of the classical slope inequality (2.16). But for $u_{f}$ large with respect to $g$ the inequality becomes interesting, as we will see in the next sections.
We now use again the algebraic index theorem in combination with Konno’s method, to obtain a somewhat different inequality as follows. Let $\Sigma$ be an irreducible horizontal component of $Z_{1}$ of maximal multiplicity $\alpha=\alpha_{1}$. The restriction of $f$ to $\Sigma$ is a (possibly ramified) covering of $\pi\colon\Sigma\to B$, of degree say $\beta$.
@@
First of all observe that $K_{f}\cdot\Sigma+\Sigma^{2}=\deg r(\pi)\geq 0$, as proved in the argument of Proposition 3.12. As $K_{f}^{2}>0$ we have that the intersection matrix between $K_{f},\Sigma$ and $F$ has determinant greater than or equal to 00:
@@
Now, plugging this inequality in (3.20), we obtain:
$K_{f}^{2}\geq(4g-4-a)\mu_{1}+K_{f}\cdot Z_{1}\geq(4g-4-a)\mu_{1}+\frac{\alpha\beta^{2}}{4(g-1)(\beta+g-1)}K_{f}^{2}.$
@@
Observe that in the case where $a=\alpha=2g-2$ (and thus $\beta=1$), we obtain (4.3). In general the inequality (4.6) seems to be weaker than (4.2), unless one has specific informations on $\beta$ and $\alpha$.
@@
Let $f\colon S\to B$ be a relatively minimal fibred surface of genus $g\geq 2$. Assume that $Z_{1}$ doesn’t have horizontal components, i.e. $a=\alpha=0$ (equivalently the linear system induced by the stalk of $\mathcal{E}_{1}$ on the general fibre is base-point-free); Then the following slope inequality holds:
@@
The inequality (4.9) can be deduced from Theorem 4.1 setting $\alpha=0$ in (4.3). ∎
@@
In the case where $f$ is a hyperelliptic fibration, Lu and Zuo in [27] proved (4.9) with the relative irregularity $q_{f}$ instead of the unitary rank $u_{f}$ In this case –again by Lu and Zuo’s results, in [28, Theorem 4.7]– there is a finite étale base change such that the flat summand of the new Hodge bundle becomes trivial, so the inequality for $q_{f}$ implies the one with $u_{f}$.
@@
The inequality (4.9) is false in general, as proved by the examples constructed by Lu in [29]. Moreover, the assumption on the base locus of the system induced by $\mathcal{E}_{1}$ is quite restrictive, and not very easy to verify in concrete cases.
Remark 4.12.
The previous theorem is somewhat close to one due to the author with Barja in [4, Remark 3.3]: there we proved that the inequality (4.9) holds if the ample summand $\mathcal{A}$ of the Hodge bundle is semistable and its stalk induces a base-point-free system on the general fibres. Recall that $\mathcal{A}$ is either the whole $f_{*}\omega_{f}$ or the one to last subsheaf of the Harder-Narasimhan filtration $\mathcal{E}_{l}$. So, the assumptions of [4, Remark 3.3] can be rephrased as follows: either the Hodge bundle is semistable (in this case by definition $Z_{1}=0$) or it has a Harder-Narasimhan filtration of length two, $\mu_{2}=0$ and $Z_{1}$ has no horizontal components. Therefore, the above Theorem implies [4, Remark 3.3].
@@
Let $f\colon S\to B$ be a relatively minimal fibred surface of genus $g\geq 2$. Let us denote as usual $r_{1}=\mbox{{rank}}\mathcal{E}_{1}$.
@@
Moreover, suppose that $f$ is locally non-trivial and that the equality is reached. Then necessarily $r_{1}=1$ and $f$ is a trigonal fibration, or $r_{1}=2$ and $g=2$. In particular, for $g\geq 3$ and $r_{1}\geq 2$, inequality (4.14) is always strict.
@@
If $u_{f}>0$, then by Remark 3.4 we have $\mathcal{A}=\mathcal{E}_{l-1}$, and the Harder-Narasimhan filtration of $\mathcal{A}$ is just the truncation of the one of $\mathcal{E}$:
@@
Let us recall the inequality (3.18):
$K_{f}^{2}\geq\left(\frac{(\alpha+1)(4g-4-a)+a}{\alpha+1}\right)\mu_{1},$
where $a=a_{1}$ and $\alpha=\alpha_{1}$. Observe that $a\geq\alpha$ and that the function
@@
If $r_{1}=1$, then $a=2g-2$, and we obtain (4.3), which coincides with the inequality (4.14).
If $r_{1}>1$, then by Clifford’s theorem applied to the subcanonical linear system ${M_{i}}_{|F}$ we have $d_{1}\geq 2(r_{1}-1)$, hence $a_{1}=2g-2-d_{1}\leq 2g-2r_{1}$. We thus have the following chain of inequalities:
$\begin{split}K_{f}^{2}&\geq\mu_{1}f(2g-2-d_{1})\geq\mu_{1}f(2g-2r_{1})=\frac{(2g-2r_{1}+1)(2g+2r_{1}-4)+2g-2r_{1}}{2g-2r_{1}+1}\mu_{1}\geq\\ &\geq\frac{(2g+2r_{1}-4)(2g-2r_{1}+1)+2g-2r_{1}}{2g-2r_{1}+1}\frac{\chi_{f}}{g-u_{f}},\end{split}$
as wanted.
Let us now suppose that the equality holds in (4.14). If $r_{1}=1$, we proved in Theorem 4.1 that necessarily $u_{f}=g-1$, and so $f$ is a trigonal fibration by Proposition 6.4 below. If $r_{2}\geq 2$, then the fibre of $\mathcal{E}_{1}$ induces a linear series on $F$, of degree $d_{1}=2r_{1}-2$. If $r_{1}<g$, again Clifford’s Theorem tells us that $F$ is hyperelliptic. So we can use the results of [27], in particular the bound of theorem 1.4, that is always strictly greater than the one of (4.14), a contradiction. If $r_{1}=g$, then the Hodge bundle is semistable, and in particular $u_{f}=0$. The equality in (4.14) becomes the equality in the classical slope inequality (2.16)
$K_{f}^{2}=\frac{4g-4}{g}\chi_{f}.$ (4.15)
We can use the inequality (4.24) of Konno, whose coefficient is strictly greater than the one of (4.15), unless $g=2$. ∎
@@
The function bounding the slope in inequality (4.14) is increasing with $u_{f}$ and $r_{1}$. For $u_{f}$ and $r_{1}$ bot small these inequalities are not meaningful. For example for the smallest cases $u_{f}=0$ and $r_{1}=1$ it is worse than the slope inequality. Notice that in the case where $r_{1}=1$, we obtain again precisely the bound (4.3) of Theorem 4.1. Note moreover that $r_{1}\leq g-u_{f}$, so the best value for $r_{1}$ is if it is precisely $g-u_{f}$. This happens under the assumption of Theorem 4.26. The strength of (4.14) lies mainly in the case where $u_{f}$ is very big.
We can somehow improve this inequality by using a sharpening of Clifford’s theorem obtained by the author with Riva in [36, Theorem 2.13]. Let us recall the result in the form needed here:
@@
Let us start two simple remarks. (1) In the case where the fibration is hyperelliptic (i.e., $c_{f}=0$), we have $m=0$ and the inequality (4.19) becomes precisely the inequality (4.14) of the theorem above. (2) If $r_{1}=1$, then the inequality (4.19) becomes the inequality (4.3).
Therefore, we have reduced ourselves to the case where $F$ is non-hyperelliptic and $r_{1}\geq 2$. In this case we can apply Theorem 4.17. Observe that
@@
Now the statement follows by applying Lemma 3.17. ∎
In the case where $r_{1}\geq 2$ we can give sharper inequalities as follows:
@@
Then one of the following possibilities holds:
@@
(i) If $h^{1}(R_{1})\geq 2$, then $R_{1}$ contributes to the Clifford index $c_{f}$ of $F$ and so $d_{1}\geq 2(h^{0}(R_{1})-1)+c_{f}\geq 2r_{1}-2+c_{f}$. So we have that $a_{1}\leq 2g-2r_{1}-c_{f}$, and we obtain equation (4.21) by using (3.18).
(ii) If $h^{1}(R_{1})=1$, in theory $R_{1}$ could be 00, but in this case we would have necessarily $d_{1}=0$ and so $r_{1}=1$, contrary to the assumption. Therefore, necessarily $R_{1}\not=0$, and by Riemann-Roch
@@
hence $a_{1}\leq g-r_{1}$, and we just use this bound again in inequality (3.18). ∎
4.3 Slope inequalities for Harder-Narasimhan filtration of length $\leq 2$
Now we prove some new slope inequalities for the case where the Harder-Narasimhan filtration is short. For the case $\ell=1$, i.e., $f_{*}\omega_{f}$ is semistable, recall the following beautiful result:
Theorem 4.23 (Konno [22]).